### v2

@@ Abstract
The evolution of cooperation has been extensively studied using abstract mathematical models and simulations. Recent advances in Large Language Models (LLM) and the rise of LLM agents have demonstrated their ability to perform social reasoning, thus providing an opportunity to test the emergence of norms in more realistic agent-based simulations with human-like reasoning using natural language. In this research, we investigate whether the cooperation dynamics presented in Boyd and Richerson’s model persist in a more realistic simulation of the diner’s dilemma using LLM agents compared to the abstract mathematical nature in the work of Boyd and Richerson. Our findings indicate that agents follow the strategies defined in the Boyd and Richerson model, and explicit punishment mechanisms drive norm emergence, reinforcing cooperative behaviour even when the agent strategy configuration varies. Our results suggest that LLM-based Multi-Agent System simulations, in fact, can replicate the evolution of cooperation predicted by the traditional mathematical models. Moreover, our simulations extend beyond the mathematical models by integrating natural language-driven reasoning and a pairwise imitation method for strategy adoption, making them a more realistic testbed for cooperative behaviour in MASs.
Keywords:
Multi-Agent Systems Large Language Model LLM Agent Social Dilemmas Agent Strategies
@@ 1 Introduction
Autonomous agents have gained significant popularity due to their immense utility in various real-world applications in customer service, healthcare, social networks, and retail domains [34]. Multi-agent systems (MASs) bring together agents with independent objectives and decision-making abilities that interact within a shared environment. As such, agents must cooperate and coordinate their actions in dynamic and often unpredictable settings [32]. Cooperative agents can help improve the performance of individual agents and the overall system [17]. Researchers have sought to understand how cooperation emerges in societies and have presented various mathematical models and simulations that predict agent behaviours [3, 6]. However, the suitability of these models for real-world, human-oriented environments remains uncertain. With the rise of alternative AI technologies, traditional rule-based decision-making approaches are being challenged, necessitating further investigation [24].
@@
Boyd and Richerson’s (B&R) model [6], a prominent simulation study of the evolution of cooperation, suggests that punishment-based mechanisms can sustain long-term cooperation. While being a simple abstract mathematical simulation, its applicability in a more realistic human-based environment remains an underexplored area. With the opportunity of using Large Language Model (LLM) agents, which demonstrate a great understanding of natural language [32], our approach consists of focusing on the diner’s dilemma. We explore different strategy compositions introduced by the B&R model to examine whether a more realistic simulation of the evolution of cooperation using LLM agents produces similar norm emergence as the abstract B&R model [6].
We model a realistic $n$-player diner’s dilemma, using LLM Agents as the backbone in making the complex dilemma decisions and allowing them to act based on their strategies, calculate payoffs for their dilemma actions, and finally reflect on their actions, analyze other agents, and change their strategies by comparing their utilities. Furthermore, following the B&R model, we allow the agent population to have multiple strategies in each simulation (currently up to four strategies per simulation). The B&R model includes experiments with both two and three strategies, but their mathematical model of population dynamics is challenging to extend beyond a small number of strategies. Furthermore, the meaning of the B&R strategies is directly encoded within their equations. In contrast, the LLM-based approach allows reasoning about strategies using natural language, making it easier to introduce and test alternative strategies in future studies. Therefore, in this study, we choose to allow different combinations of four strategies in the population, making the simulation more complex and realistic, enabling a deeper analysis of the strategy evolution.
In summary, we aim to investigate the impact of strategies from B&R’s work and their evolution with repeated diner’s dilemma scenario, modelled with novel LLM agents, which have been shown to implicitly capture human reasoning and thinking abilities, thus, allowing us to gain insights on whether these LLM Agents behave similarly as shown in the abstract mathematical simulation studies.
@@ 2.1 Game Theory and Social Dilemmas
Social dilemmas are an agent-related strategic concept that motivates the study of the normative behaviour of agents in MASs. A social dilemma occurs when an agent is forced to choose between actions that maximize their personal gain at the expense of the group’s collective benefit or actions that promote the collective good but lower their personal benefits [16]. This scenario is essentially a conflict between personal and social optimality. As a result, agents who aim to satisfy their short-term self-interests are often characterized as non-cooperative as they are less likely to choose actions that serve the long-term benefit of the group. These social dilemmas can be explained through game theory, which analyzes how rational agents make decisions when faced with interactions with individuals in competitive or cooperative game environments. Game theory helps identify vital insights in explaining the behaviour of agents under economic, political and social interaction [7, 4].
A dilemma requires the agent or individual to decide whether to cooperate with or to defect against its opponent. Based on these decisions, four main payoff values are defined, as elaborated by Macy and Flache [21]: (i) reward (R), given when both agents choose to cooperate, (ii) punishment (P), incurred when both agents defect, (iii) temptation (T), where agent defects and unfairly benefits while the opponent cooperates, and (iv) sucker cost (S), where a cooperating agent suffers a loss when the opponent defects. These payoff values form the foundation of various social dilemmas studied in game theory, such as the Prisoner’s Dilemma, the chicken game, stag hunt [34] and the trust game [20]. These two-player social dilemmas can be redesigned into their $n$-player form [14, 20], which closely relates to real-life examples [22]. In the Diner’s Dilemma [36], a group of agents agrees to split the cost of their meals. However, individual agents may exploit this arrangement by ordering expensive items and transferring a portion of their costs onto the group.
@@ 2.3 LLM Agents
Despite extensive research on the effects of metanorms and punishment mechanisms to induce cooperation, limited work has been conducted in conjunction with language models. In the few works applying LLMs to reasoning about social dilemmas, LLMs struggle with such interactions—GPT-4111https://openai.com/index/gpt-4 has been shown to select actions that maximize its personal gain and fails to coordinate with fellow agents in games such as the Battle of the Sexes [2] and frequently selects uncooperative actions that harshly penalize minor mistakes by opponents. Therefore, although LLMs exhibit strong alignment with human behaviour, they struggle to achieve the high levels of cooperation seen in real-world human interactions. This limitation suggests that LLMs should be carefully evaluated when integrated into social experiments [9].
@@ 2.4 Agent Simulations
Simulations are commonly utilised to explore AI agent behaviour within a virtual environment [38]. These frameworks use research from multiple fields, such as social sciences, psychology, economics and AI for understanding social phenomena. To study social behaviour in group settings, it is vital to develop simulations that closely depict human activities and track changes in the world states, such as movements of objects resulting from agent actions [40]. These simulations are broadly categorized as task-based or social interaction-based simulations. Task-based simulations [11], such as the ScienceWorld environment, are used for conducting science experiments using a text-based framework [39, 15], whereas social simulations, such as Melting Pot [1], use multi-agent reinforcement learning environments. The use of LLM agents in simulations provides a more realistic approach in replicating normative behaviours [19, 32], as seen in frameworks like AgentVerse [8], but with representation of the objectives in an abstract manner rather than using specific social scenarios. Game engines like ALFWorld [33] and Watch-And-Help [29] provide detailed environmental control but are often complex to modify and lack seamless LLM agent integration. Meanwhile, a sandbox environment, Smallville [^5] https://github.com/nickm980/smallville, provides an interactive and customizable LLM agent-driven simulation, which has been utilized to build agent societies [28] and also normative frameworks [31]. Thus, we consider that this provides a promising simulation framework for experimenting with the work by B&R in a realistic manner.
@@
In summary, to the best of our knowledge, no papers have studied the evolution of cooperation with the strategies introduced in the B&R work using LLM agents. Although some articles have explored dilemmas, such as the Prisoner’s Dilemma with LLMs as well as other strategies, there is a lack of research on exploring the evolutionary aspect of the cooperation strategies within a more realistic, $n$-player dilemma scenario to infer insights on whether it produces similar norm emergence as in the abstract B&R work.
@@ 3.1 Simulation Environment Setup
The simulation framework was adapted from the existing Smallville environment utilized in the CRSEC framework [31] [^6] https://github.com/sxswz213/CRSEC as it is a promising framework for experimenting with the work of B&R in a realistic manner, as explained in Section 2.4. The framework can be utilized to model realistic social dilemmas, specifically, The Diner’s Dilemma, as illustrated in Figure 1. The virtual environment was designed using the Tiled Map Editor [^7] https://www.mapeditor.org/ for layout and Phasor [^8] https://phaser.io/ for agent movement, creating two primary settings: a pub and a cafe, where the agents interact. A total of eight agents were introduced and divided into two groups, where each agent was assigned distinct strategies and lifestyles to simulate diverse attributes and interactions. Agent lifestyles—a concept used in the works of Ren et al. [31]—can be used to simulate human-like behaviour (e.g., “Likes to take a high-intensity run in the morning and needs high nutrition for it”) in our LLM agents and to test for biases in LLM decision prompting when engaging in social dilemmas.
@@
In our research, we encode these strategies as part of the prompts provided to the LLM, guiding the agent decision-making process in social dilemma situations [18]. As agents navigate various scenarios in the simulated environment, norms emerge when a significant portion of the population adopts successful strategies transmitted through evolutionary mechanisms. Hence, our objective is to investigate the dynamics that emerge from the interactions of these strategies within the population of the simulated environment.
To ensure scalability and adaptability, our simulation system calls the Groq API [^9] https://console.groq.com/docs/api-reference and Ollama [^10] https://ollama.com/, leveraging the latest advancements in large language models (LLMs).
@@ 3.2 Diner’s Dilemma Simulation Process
The simulation of a Diner’s Dilemma involved multiple stages, as illustrated in Figure 2, each incorporating LLM-based decision-making processes. In this scenario, a group of agents meet in the cafe or the pub, having agreed to split the cost of their meals. Each agent faces a dilemma in deciding what type of meal to order from the given two options: budget or premium. The agent can either cooperate by choosing the less preferred budget option to increase the collective benefit or choosing the preferred premium (and more expensive) option to maximize the personal gain. Then, the agents apply their individual strategies to determine whether to punish defectors (defined as agents who choose the premium meal) and, in the case of Moralists, to punish those who fail to punish defectors, thereby enforcing metanorms. Next, the agents update their individual utilities according to their decisions in ordering and considering the punishment costs. Finally, the agents compare their utilities with other agents and determine whether to adopt a different strategy. The individual stages are described in detail below.
@@
    Punishment Stage (Evaluate Cooperation - Round 1): Agents evaluate the actions of others and decide whether to punish defectors. LLMs are used to decide whether to scold one another, mainly based on the agent’s strategy. Other inputs consist of the order history (Figure 4).
@@
    Metanorm Enforcement Stage (Evaluate Cooperation - Round 2): Moralists punish not only defectors but also non-punishers of defectors. This stage involves higher-order normative reasoning (i.e., a metanorm), where moralist agents punish those who have failed to enforce norms (Figure 5).
@@
    Utility Assessment and Strategy Update: Each agent’s actions and outcomes are logged to update a numerical utility score that reflects its current performance. Here, apart from the order history, punishment costs are also considered when updating the utility. Punishing another agent will incur a cost of $k$ to the punisher and a cost of $p$ to the agent who is being punished (where usually $p$ > $k$). To determine whether an agent should adopt a new strategy, we utilize a pairwise imitation method based on the Fermi function [37] to drive the spread of successful strategies. Here, with the payoffs calculated during the diner’s dilemma, the utility of the focal agent (A) is compared against that of a randomly chosen role model (B) using the following equation from the Fermi process, which computes the probability of agent A changing to use the target agent’s strategy.
@@ 4 Experimentation
Our experiments were conducted using the open-source model Meta-Llama-3-70B-Instruct provided through the Groq API without any fine-tuning; we used the original model as provided. We conducted separate testing for each of the four stages where the LLM is used by varying the temperature and top_p (or nucleus sampling, where tokens with top_p probability mass are considered), different strategies, and lifestyles of personas (e.g., Morning Runner, Newspaper Reader, Photographer) to assess biases and to ensure expected results are achieved. Finally, we integrated all of the tested prompts and LLM agents with strategies and lifestyles into the simulation.
@@
The first group of the first combination represents a balanced population with all the strategies present. We chose double moralists for the second group of the same combination, especially with a P and E being present in this combination (groups 1 and 2 together), to see the metanorm punishments in effect. Even though P punishes the defector R1, E will not punish; therefore, the moralist should interfere and punish both E and P because P is failing to punish a non-punisher (E) as well. The first group of the second combination was chosen to see the effect when initialized with more R1 agents in the population, whereas group 2 was chosen to represent a population with more punishers (P). Therefore, our selection covers both a balanced group and groups where a majority of agents share the same strategy. In this initial study, we have focused on this set of strategy combinations to investigate the cooperation dynamics from B&R’s model. Figure 8 illustrates an example scenario in the simulation of the agent interactions in different stages in the diner’s dilemma.
@@
Our selection for the punishment costs for p and k ($k$, the cost of punishment, and $p$, the cost of being punished) is based on the equations presented in the B&R model [6] and in [5] where for small populations, we can vary p by fixing k to observe the change in the population dynamics. Hence, whenever we define the punishment costs values to the LLM agents, we set k=1 while increasing p.

### v3

@@ Abstract
The evolution of cooperation has been extensively studied using abstract mathematical models and simulations. Recent advances in Large Language Models (LLMs) and the rise of LLM agents have demonstrated their ability to perform social reasoning, thus providing an opportunity to test the emergence of norms in more realistic agent-based simulations with human-like reasoning using natural language. In this research, we investigate whether the cooperation dynamics presented in Boyd and Richerson’s model persist in a more realistic simulation of the Diner’s Dilemma using LLM agents compared to the abstract mathematical nature in the work of Boyd and Richerson. Our findings indicate that agents follow the strategies defined in the Boyd and Richerson model, and explicit punishment mechanisms drive norm emergence, reinforcing cooperative behaviour even when the agent strategy configuration varies. Our results suggest that LLM-based Multi-Agent System simulations, in fact, can replicate the evolution of cooperation predicted by the traditional mathematical models. Moreover, our simulations extend beyond the mathematical models by integrating natural language-driven reasoning and a pairwise imitation method for strategy adoption, making them a more realistic testbed for cooperative behaviour in MASs.
@@ 1 Introduction
Autonomous agents have gained significant popularity due to their immense utility in various real-world applications in customer service, healthcare, social networks, and retail domains [35]. Autonomous agents can even co-exist with humans in virtual environments [31]. Multi-agent systems (MASs) bring together agents with independent objectives and decision-making abilities that interact within a shared environment. As such, agents must cooperate and coordinate their actions in dynamic and often unpredictable settings [33]. Cooperative agents can help improve the performance of individual agents and the overall system [17]. Researchers have sought to understand how cooperation emerges in societies and have presented various mathematical models and simulations that predict agent behaviours [3, 6]. However, the suitability of these models for real-world, human-oriented environments remains uncertain. With the rise of generative AI technologies, traditional rule-based decision-making approaches are being challenged, necessitating further investigation [24].
@@
Boyd and Richerson’s (B&R) model [6], a prominent simulation study of the evolution of cooperation, suggests that punishment-based mechanisms can sustain long-term cooperation. While being a simple abstract mathematical simulation, its applicability in a more realistic human-based environment remains an underexplored area. With the opportunity of using Large Language Model (LLM) agents, a form of generative AI technology, which demonstrate a great understanding of natural language [33], we explore different strategy compositions introduced by the B&R model to examine whether a more realistic simulation of the evolution of cooperation using LLM agents produces similar norm emergence as the abstract B&R model [6].
We model a realistic $n$-player Diner’s Dilemma, using LLM agents as the backbone in making the complex dilemma decisions and allowing them to act based on their strategies, calculate payoffs for their dilemma actions, and finally reflect on their actions, analyze other agents, and change their strategies by comparing their utilities. Furthermore, following the B&R model, we allow the agent population to have multiple strategies in each simulation (currently up to four strategies per simulation). The B&R model includes experiments with both two and three strategies, but their mathematical model of population dynamics is challenging to extend beyond a small number of strategies. Furthermore, the meaning of the B&R strategies is directly encoded within their equations. In contrast, our LLM-based approach allows reasoning about strategies using natural language, making it easier to introduce and test alternative strategies in future studies. Therefore, in this study, we choose to allow different combinations of four strategies in the population, making the simulation more complex and realistic, enabling a deeper analysis of the strategy evolution.
In summary, we aim to investigate the impact of strategies from B&R’s work and their evolution with repeated Diner’s Dilemma scenario, modelled with novel LLM agents, which have been shown to implicitly capture human reasoning and thinking abilities, thus allowing us to gain insights into whether these LLM agents behave similarly as shown in the abstract mathematical simulation studies. Our findings suggest that LLM agents in a Diner’s Dilemma simulation show promising convergence toward cooperative strategies under explicit increasing punishment costs.
@@ 2.1 Game Theory and Social Dilemmas
Social dilemmas are an agent-related strategic concept that motivates the study of the normative behaviour of agents in MASs. A social dilemma occurs when an agent is forced to choose between actions that maximize their personal gain at the expense of the group’s collective benefit or actions that promote the collective good but lower their personal benefits [16]. This scenario is essentially a conflict between personal and social optimality. As a result, agents who aim to satisfy their short-term self-interests are often characterized as non-cooperative, as they are less likely to choose actions that serve the long-term benefit of the group. These social dilemmas can be explained through game theory, which analyzes how rational agents make decisions when faced with interactions with individuals in competitive or cooperative game environments. Game theory helps identify vital insights in explaining the behaviour of agents under economic, political and social interaction [7, 4].
A dilemma requires the agent or individual to decide whether to cooperate with or to defect against its opponent. Based on these decisions, four main payoff values are defined, as elaborated by Macy and Flache [21]: (i) reward (R), given when both agents choose to cooperate, (ii) punishment (P), incurred when both agents defect, (iii) temptation (T), where agent defects and unfairly benefits while the opponent cooperates, and (iv) sucker cost (S), where a cooperating agent suffers a loss when the opponent defects. These payoff values form the foundation of various social dilemmas studied in game theory, such as the Prisoner’s Dilemma [18, 14], the chicken game [21], stag hunt [35] and the trust game [20]. These two-player social dilemmas can be generalised into their $n$-player form [14, 20]. In the Diner’s Dilemma [37], a group of agents agrees to split the cost of their meals. However, individual agents may exploit this arrangement by ordering expensive items and transferring a portion of their costs onto the group.
@@ 2.3 LLM Agents
Despite extensive research on the effects of metanorms and punishment mechanisms to induce cooperation, limited work has been conducted in conjunction with LLMs. In the few works applying LLMs to reasoning about social dilemmas, LLMs struggle with such interactions—GPT-4111https://openai.com/index/gpt-4 has been shown to select actions that maximize its personal gain and fails to coordinate with fellow agents in games such as the Battle of the Sexes [2] and frequently selects uncooperative actions that harshly penalize minor mistakes by opponents. Therefore, although LLMs exhibit strong alignment with human behaviour, they struggle to achieve the high levels of cooperation seen in real-world human interactions. This limitation suggests that LLMs should be carefully evaluated when integrated into social experiments [9].
@@ 2.4 Agent Simulations
Simulations are commonly utilised to explore AI agent behaviour within a virtual environment [39]. These frameworks use research from multiple fields, such as social sciences, psychology, economics and AI for understanding social phenomena. To study social behaviour in group settings, it is useful to develop simulations that closely depict human activities and track changes in the world states, such as movements of objects resulting from agent actions [41]. These simulations are broadly categorized as task-based or social interaction-based simulations. Task-based simulations [11], such as the ScienceWorld environment, are used for conducting science experiments using a text-based framework [40, 15], whereas social simulations, such as Melting Pot [1], use multi-agent reinforcement learning environments. The use of LLM agents in simulations provides a more realistic approach in replicating normative behaviours [19, 33], as seen in frameworks like AgentVerse [8], but with representation of the objectives in an abstract manner rather than using specific social scenarios. Game engines such as ALFWorld [34] and Watch-And-Help [29] provide detailed environmental control but are often complex to modify and lack seamless LLM agent integration. Meanwhile, a sandbox environment, Smallville [^5] https://github.com/nickm980/smallville, provides an interactive and customizable LLM agent-driven simulation, which has been utilized to build agent societies [28] and also normative frameworks [32]. Thus, we consider that this provides a promising simulation framework for experimenting with the work by B&R in a realistic manner.
@@
In summary, to the best of our knowledge, no research has studied the evolution of cooperation with the strategies introduced in the B&R work using LLM agents. Although some articles have explored dilemmas, such as the Prisoner’s Dilemma with LLMs as well as other strategies, there is a lack of research on exploring the evolutionary aspect of the cooperation strategies within a more realistic, $n$-player dilemma scenario to infer insights on whether it produces similar norm emergence as in the abstract B&R work.
@@ 3.1 Simulation Environment Setup
The simulation framework was adapted from the existing Smallville environment utilized in the CRSEC framework [32] [^6] https://github.com/sxswz213/CRSEC as it is a promising framework for experimenting with the work of B&R in a realistic manner, as explained in Section 2.4. The framework was utilized to model realistic social dilemmas, specifically, The Diner’s Dilemma, as illustrated in Figure 1. The virtual environment was designed using the Tiled Map Editor [^7] https://www.mapeditor.org/ for layout and Phasor [^8] https://phaser.io/ for agent movement, creating two primary settings: a pub and a cafe, where the agents interact. A total of eight agents were introduced and divided into two groups, where each agent was assigned distinct strategies and lifestyles to simulate diverse attributes and interactions. Agent lifestyles—a concept used in the works of Ren et al. [32]—can be used to simulate human-like behaviour (e.g., “Likes to take a high-intensity run in the morning and needs high nutrition for it”) in our LLM agents and to test for biases in LLM decision prompting when engaging in social dilemmas.
@@
In our research, we encode these strategies as part of the prompts provided to the LLM, guiding the agent’s decision-making process in social dilemma situations [18]. As agents navigate the scenario in the simulated environment, we consider a norm to have emerged when a significant proportion of the population adopts a successful strategy transmitted through evolutionary mechanisms. Hence, our objective is to investigate the dynamics that emerge from the interactions of these strategies within the population of the simulated environment.
@@ 3.2 Diner’s Dilemma Simulation Process
The simulation of a Diner’s Dilemma involved multiple stages, as illustrated in Figure 2, each incorporating LLM-based decision-making processes. In this scenario, a group of agents meet in the cafe or the pub, having agreed to split the cost of their meals. Each agent faces a dilemma in deciding the type of meal to order from the given two options: budget or premium. The agent can either cooperate by choosing the less preferred budget option to increase the collective benefit or choosing the preferred premium (and more expensive) option to maximize the personal gain. Then, the agents apply their individual strategies to determine whether to punish defectors (defined as agents who choose the premium meal) and, in the case of Moralists, to punish those who fail to punish defectors, thereby enforcing metanorms. Next, the agents update their individual utilities according to their decisions in ordering and considering the punishment costs. Finally, the agents compare their utilities with other agents and determine whether to adopt a different strategy. The individual stages are described in detail below.
@@
    Punishment Stage (Evaluate Cooperation - Round 1): Agents evaluate the actions of others and decide whether to punish defectors. LLMs are used to decide whether to scold one another based on the agent’s strategy. Other inputs consist of the order history (Figure 4).
@@
    Metanorm Enforcement Stage (Evaluate Cooperation - Round 2): This stage involves higher-order normative reasoning (i.e., a metanorm), where moralist agents punish those who have failed to enforce norms (Figure 5).
@@
    Utility Assessment and Strategy Update: Each agent’s actions and outcomes are logged to update a numerical utility score that reflects its current performance. Here, apart from the order history, punishment costs are also considered when updating the utility. Punishing another agent will incur a cost of $k$ to the punisher and a cost of $p$ to the agent who is being punished (where usually $p$ > $k$). In small populations, changes in population dynamics can be observed by varying p while keeping k fixed [6, 5]. Accordingly, when assigning punishment cost values to the LLM agents, we set k = 1 and increased the value of p. To determine whether an agent should adopt a new strategy, we utilize a pairwise imitation method based on the Fermi function [38] to drive the spread of successful strategies. Here, with the payoffs calculated during the Diner’s Dilemma, the utility of the focal agent (A) is compared against that of a randomly chosen role model (B) using the following equation from the Fermi process, which computes the probability of agent A changing to use the target agent’s strategy.
@@ 4 Experimentation
Our experiments were conducted using the open-source model Meta-Llama-3.3-70b-Instruct provided through the Vertex API by Google Cloud without any fine-tuning; we used the original model as provided. We conducted separate testing for each of the four stages where the LLM is used by varying the temperature (randomness of text that is generated by the LLM) and top_p (or nucleus sampling, where tokens with top_p probability mass are considered), different strategies, and lifestyles of personas (e.g., Morning Runner, Newspaper Reader, Photographer) to assess biases and to ensure expected results are achieved. Finally, we integrated all of the tested prompts and LLM agents with strategies and lifestyles into the simulation.
@@
The first group of the first combination represents a balanced population with all the strategies present. We chose double moralists for the second group of the same combination, especially with a P and E being present in this combination (groups 1 and 2 together), to see the metanorm punishments in effect. Even though P punishes the defector R1, E will not punish; therefore, the moralist should interfere and punish both E and P because P is failing to punish a non-punisher (E) as well. The first group of the second combination was chosen to see the effect when initialized with more R1 agents in the population, whereas group 2 was chosen to represent a population with more punishers (P). Therefore, our selection covers both a balanced group and groups where a majority of agents share the same strategy. In this initial study, we have focused on this set of strategy combinations to investigate the cooperation dynamics from B&R’s model. Figure 8 illustrates an example scenario in the simulation of the agent interactions in different stages of the Diner’s Dilemma.
@@
The two agent behaviour distributions were chosen to represent contrasting environments: one with a higher proportion of M agents (Combination 1) and another dominated by R1 agents (Combination 2). To examine the impact of enforcing punishment within these settings, we considered three variations of the punishment parameters $p$ and $k$. Due to computational challenges in scaling to a larger number of iterations, only a limited set of experiments could be carried out in this preliminary study. This is intended to be addressed in future works through a thorough systematic testing process.